OpenStudy (anonymous):

Determine the range of f if f(x)= 2x^2+5x-3

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

how do u write that in interval notation

OpenStudy (anonymous):

[-inf,inf]

OpenStudy (anonymous):

says tahts wrong

OpenStudy (anonymous):

thats the domain

OpenStudy (anonymous):

yes thats the domain - sorry whats the range then - ill check that out

OpenStudy (anonymous):

yea i need the range please. I dont know it

OpenStudy (anonymous):

its x vertex : -b/2a = -5/4 so its y vertex : 2*(-5/4)^2 + 5*(-5/4) -3 = -6.125 since a (2) is positive the parabola is concave upwards .. so -6.125 is the smallest y possible.. (minimum point) [-6.125,infinity]

OpenStudy (anonymous):

i think it may be wrong

OpenStudy (anonymous):

[-6.125,infinity)

OpenStudy (anonymous):

ya its still wrong i think

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

idk then it gives me lots of questions in one and wont tell me whch one is wrong

OpenStudy (anonymous):

post the questions